OpenStudy (caitlinnr14):

Help PLEASE! ]Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

y=x/7-3 so y intercept would equal (0,-3)

OpenStudy (caitlinnr14):

so wait how did u get tht

OpenStudy (caitlinnr14):

oh nvm i see u divided 7 and 21

OpenStudy (caitlinnr14):

-21

OpenStudy (anonymous):

yupp

OpenStudy (caitlinnr14):

so it would be the last one

OpenStudy (anonymous):

let me check the x intercept first i just knew the y right off the top of my head

OpenStudy (anonymous):

make R.H.S =1 by dividing through out by -21 \[\frac{ x }{-21 }+\frac{ 7y }{-21 }=1\] \[or \frac{ x }{ -21 }+\frac{ y }{-3 }=1\] x intercept =-21,y intercept=-3

OpenStudy (anonymous):

kool ive never seen it done like that before

OpenStudy (caitlinnr14):

i think it would be the lasst one

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

it is equation of line in intercept form i.e., \[\frac{ x }{ a }+\frac{ y }{ b }=1\]

OpenStudy (caitlinnr14):

ok thx @iLuis21 and @surjithayer

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

yw